# MnLargeSymbols’164 MnLargeSymbols’171

### v2

@@ II.1 The Extended Yard Sale model
In the Extended Yard Sale model (EYS) [11], NN agents exchange wealth following a stochastic process where $w_{i}(t)$ denotes the wealth of agent $i$ at time $t$. The total wealth is conserved, so $W=\sum_{i}w_{i}(t)$ for all $t$. At each time step $t\to t+\Delta t$, two agents $i$ and $j$ are chosen at random to exchange wealth and pay or receive benefits from the tax system according to the rule
@@
for agent $i$ and an analogous rule for agent $j$. The second term of Eq. II.1 represents the social benefit earned minus the tax payed by agent ii. Its action redistributes wealth from the rich ($w_{i}>W/N$) to the poor ($w_{i}<W/N$) agents. The third term represents the amount of wealth exchanged between $i$ and $j$ in the interaction. By definition of the YS model, it is a fraction of the wealth of the poorer agent, so
@@ II.2 Networks
The structure of networks is usually represented by graphs, which in turn are composed by a set of nodes and a set of links. The number of directed links of source $j$ and target $i$ of a network of NN nodes in $\{1,...,N\}$ can be represented by the entry $a_{ij}\in\{0,1,2,$...\} of a corresponding adjacency matrix $a$ [36]. The number of outgoing and incoming links of node $i$ are denoted by $k_{i}=\sum_{j}a_{ij}$ and $q_{i}=\sum_{j}a_{ji}$ respectively, $M=\sum_{ij}a_{ij}=\sum_{i}k_{i}=\sum_{j}q_{j}=N\bar{k}=N\bar{q}$ denotes the number of directed links, and $\bar{k}$ and $\bar{q}$ the average out- and in-degree of the nodes of the network  [^1] Note that, usually, $2M=\sum_{ij}a_{ij}$ is used for nondirected networks. Here $M=\sum_{ij}a_{ij}$ is preferred since it works for both, directed and non-directed networks.. Non-directed networks are represented with symmetric adjacency matrices, in which case the number of non-directed links is $M/2$, $k_{i}=q_{i}$ and, by convention, $a_{ii}\in\{0,2,4,$...\} for each ii.
@@ II.5 Interaction and taxation modes over networks
Different taxation and interaction modes can be considered. Each interaction-taxation event is represented by a tuple $s=(uv,l,m)$ where the pair uvuv denotes a directed link chosen from a network $a$ representing an interaction initiated by an agent $u$ over an agent $v$, and $l$ and $m$ represent the pair of agents that are being taxed. Without loosing generality, at each event two agents are taxed simultaneously to unify different taxation modes within the same representation. The outcomes of uvuv, $l$ and $m$ affect the statistics of $b$ and $c$. Specifically, $P(c_{ij}=1|s,a)=P(c_{ij}=1|uv)=\delta_{ij,uv}+\delta_{ij,vu}$, $P(c_{ij}=0|s,a)=1-P(c_{ij}=1|s,a)$, $P(b_{i}=1|s,a)=P(b_{i}=1|l,m)=\delta_{il}+\delta_{im}$ and $P(b_{i}=0|s,a)=1-P(b_{i}=1|s,a)$.
Two interaction modes are considered. In Interaction Mode A (IMA), a directed link uvuv is chosen uniformly at random among the $M$ available. This occurs with probability
@@
For each interaction mode, two taxation modes are also considered. In Taxation Mode A (TMA), the agents $u$ and $v$ participating in the interaction uvuv are the ones being taxed. Formally, $l=u$ and $m=v$, so $P(l,m|uv)=P(l|uv)P(m|uv)=\delta_{lu}\delta_{mv}$. Therefore
@@
where $h_{i}=\sum_{u}a_{ui}/k_{u}\approx\sum_{u}a_{ui}/\bar{k}=q_{i}/\bar{k}$ if $1/k_{u}\approx 1/\bar{k}$ is assumed for all $u$ in the neighborhood of ii. In Taxation Mode B, the pair of agents being taxed are uniformly chosen at random. Hence, $P(l,m|uv)=P(l|uv)P(m|uv)=N^{-2}$ and, therefore
@@ III Results
Figure 1: Behavior of EYS model on fully connected networks. a) The expected relative wealth $x_{1}$ owned by the richest agent at the stationary regime is plotted vs temperature $T$ for different system sizes NN, averaged over 64 MC simulations of the stochastic process of Eq. II.1 starting from uniform initial conditions. b) A comparison of $x_{1}$ vs $T$ for $N=128$, showing results from MC simulations (blue circles), a single numerical simulation of the QMF approximation of Eq. II.4 starting from a uniform initial condition with added Gaussian noise (mean = 0, standard deviation = 0.01, green squares), and the MF prediction of Eq. 13 (red solid line). c) In a system of $N=16$ agents, each line represents the time evolution of the $w_{i}/W$ for i=1,,16i=1,...,16 in a MC simulation below the critical temperature, $T=0.25$. The inset shows analogous results but above the critical temperature, $T=1.1$. d) Time evolution of the same system as in c) but for a single numerical simulation of the QMF approximation at $T=0.25$ (main plot) and $T=1.1$ (inset).
@@ III.1 Fully connected networks
Figure 1a) shows results for MC simulations of the stochastic process of Eq. II.1 for different sizes NN, which corresponds to the EYS model on fully connected networks. Except when stated otherwise, the uniform state $x_{i}=1/N$ is used as the initial condition for all simulations. The expected relative wealth of the richest agent $x_{1}$ in the steady state is plotted vs the temperature $T$. The phase transition is confirmed and finite size effects are observed. In Fig. 1b) the MC simulations are compared against numerical predictions of the QMF approximation of Eq. II.4 and the analytical MF predictions derived from Eq. III.1. Here, a small Gaussian noise is added to the initial condition to break the symmetry. As can be seen, the predictions agree up to finite size effects. Examples of the microscopic dynamics from simulations of the MC and the QMF approximations of a system with $N=16$ agents are shown in Figs. 1c) and 1d), respectively. Moreover, the wealth of multiple agents increases with time during an initial transient period. Note, however, this growth is short-lived. Eventually, all agents with the exception of the richest begin to experience a decline in wealth. This behavior arises because wealthier agents accumulate resources at the expense of their neighbors. During the early stages, several secondary wealthy agents see their wealth grow while they extract resources from poorer neighbors. Once these surrounding resources are exhausted, only the competition between the rich agents remain, ultimately favoring only the wealthiest.
@@ III.2 Erdős-Rényi random networks
The numerical experiments show that, for all interaction and taxation modes, the value of $\left\langle x_{1}\right\rangle_{a}$ grows with $\bar{k}$ but decays with NN at zero or small temperatures, $T\gtrsim 0$ (not shown). In other words, there is a competition between $\bar{k}$ and NN for consolidating or weakening the condensed phase, respectively. To better characterize this behavior, assume the scaling hypothesis
@@
for arbitrary values of $\lambda$, $\bar{k}$, NN and zero or small temperatures. Choosing the value of $\lambda$ such that $\lambda^{\mu}N=1$, the scaling relation
@@
is obtained, for some universal function $g_{T}$ that generally depends on $T$. As shown in Fig. 3, a collapse of the curves $N^{-1/\mu}\left\langle x_{1}\right\rangle_{a}(\bar{k},N)$ vs $N^{-\nu/\mu}\bar{k}$ is obtained for different system sizes NN while varying $\bar{k}$. By visual inspection, it is found that the scaling occurs for $\nu=\mu=\infty$ such that $\nu/\mu=1$ when $T=0$ (Figs. 3 a) and b)) and for $\nu=\mu=-4$ when $T=0.06$ (Figs. 3 c) and d)). In other words, the scaling
@@
Figure 3: Scaling of the network average of the expected relative wealth of the richest agent $x_{1}$ at the stationary regime, with respect to network sizes NN and average degrees $\bar{k}$, for MC simulations (left column) and QMF approximations (right column), at two specific temperatures, $T=0$ (top row) and $T=0.06$ (bottom row). In all cases, the stars are obtained at $\bar{k}=N-1$, i.e. the fully connected case. In panel d), the dashed cyan curve corresponds to the scaling $g_{T}(z)=(1-z)^{-1/4}$ for $z=(N-1)/N\lesssim 1$ and the red solid curve to the scaling $g_{T}(p)=-1/3+p^{1/3}(1-p)^{-1/4}$ for $p\approx\bar{k}/N$. In the inset, the same results are shown in log-log scale.
@@
which captures the imbalance between poor and rich agents. By construction, $\phi\to 0$ as $T\to\infty$, and $\phi\approx 1$ in the low-temperature and high-connectivity limit $T\to 0$, $\bar{k}\to N-1$. These predictions for $\phi$ are corroborated by both MC simulations and QMF approximations, as shown in Figs. 5 a) and b). In particular, Fig. 5a) reveals a crossover at a characteristic temperature $T=T_{c}$, which appears consistently across different values of $\bar{k}$. Moreover, while the transition is continuous in MC simulations, it becomes discontinuous in the QMF approximation as $\bar{k}\to N-1$, likely due to the mean-field treatment. The behavior of the time-averaged order parameter $\langle\phi\rangle_{a}$ also reflects the interplay between connectivity $\bar{k}$ and system size NN. For $T<T_{c}$, $\langle\phi\rangle_{a}$ increases with $\bar{k}$ and decreases slightly with NN; the opposite trend is observed for $T>T_{c}$. These trends are consistent with the approximation $r_{\max}\approx(2/\bar{k})\ln N$ derived from Eq. 21 for $\bar{k}\ll N$ and $T\approx 0$, which predicts a logarithmic growth with NN and an inverse dependence on $\bar{k}$.
@@
In Fig. 5d) network averages of the cumulative wealth of the rich agents, $u^{*}R^{*}$ are plotted as a function of $T$ for different network sizes NN and a fixed link density $p\approx\bar{k}/N$ for QMF approximations (main plot) and MC simulations (inset). A clear change in behavior is observed just below $T_{c}$, in agreement with the existence of a phase transition. In particular, for QMF approximations, there is a non-monotonous behavior where the minimum of $u^{*}R^{*}$ drops with NN because $R^{*}$ decays faster than the growth of $u^{*}$ at such point. This behavior is not observed in MC where $R^{*}$ tends to overestimate the effective number of rich agents due to stochastic fluctuations.
@@
Figure 5: a) The steady state network average of the order parameter $\phi$ defined in Eq. 22 is plotted as a function of temperature $T$ for networks of $N=128$ nodes and different average degrees $\bar{k}$, using IMA and TMA, for QMF approximations. In the inset, analogous results for for MC simulations. b) Similar to panel a) and corresponding inset, but fixing the average degree $\bar{k}=8$ while varying the network size NN. c) The network average of the average wealth $u^{*}$ of the rich agents is plotted as a function of $T$ by direct calculation with MC simulations (orange circles) and QMF approximations (purple symbols) for $N=128$ and $\bar{k}=8$. For comparison, corresponding predictions from Eq. 19 obtained by direct calculations of network averages of the number $R^{*}$ of rich agents are included for MC simulations and QMF approximations (solid lines of respective colors). d) The network average of the aggregated wealth of the rich agents $R^{*}u^{*}$ is plotted as a function of the temperature $T$ for different values of NN, keeping fixed the link density $\bar{k}/N=0.5$, for QMF approximations (main plot) and MC simulations (inset).
@@
In Fig. 6 the network average of the distribution of normalized wealths $P(x)$ is shown for QMF approximations in different scenarios. The top row shows distributions for all interaction and transaction modes for relatively sparse networks. In the middle and bottom row, distributions for the IMA/TMA modes are shown for varying degrees $\bar{k}$ and network sizes NN, respectively. In the left column, the distribution is shown for the condensed phase. In all cases two modes or peaks are observed, one for poor agents and another for rich agents (Figs. 6 a)d) and g)). This result has also been observed in the MC simulations. In the middle column, distributions obtained for temperatures near or below the critical point are shown. Broad distribution are observed, except for the IMB/TMA case, where there is a sharp cutoff at the tail of the curve (Fig. 6b)). The distribution broadens as the average degree $\bar{k}$ decreases for IMA/TMA (Fig. 6e)). Interestingly, this effect is also observed at high temperatures (Fig. 6f)). However, as seen in Fig. 6c), this is not the case for the other modes, especially for TMB where the distribution is highly concentrated around $x=1/N$. Finally, as shown in Figs. 6 g)h) and i), the distributions suffer little variation with network size.

### v3

@@ II.1 The Extended Yard Sale model
In the Extended Yard Sale model (EYS) [11], $N$ agents exchange wealth following a stochastic process where $w_{i}(t)$ denotes the wealth of agent $i$ at time $t$. The total wealth is conserved, so $W=\sum_{i}w_{i}(t)$ for all $t$. At each time step $t\to t+\Delta t$, two agents $i$ and $j$ are chosen at random to exchange wealth and pay or receive benefits from the tax system according to the rule
@@
for agent $i$ and an analogous rule for agent $j$. The second term of Eq. II.1 represents the social benefit earned minus the tax payed by agent $i$. Its action redistributes wealth from the rich ($w_{i}>W/N$) to the poor ($w_{i}<W/N$) agents. The third term represents the amount of wealth exchanged between $i$ and $j$ in the interaction. By definition of the YS model, it is a fraction of the wealth of the poorer agent, so
@@ II.2 Networks
The structure of networks is usually represented by graphs, which in turn are composed by a set of nodes and a set of links. The number of directed links of source $j$ and target $i$ of a network of $N$ nodes in $\{1,...,N\}$ can be represented by the entry $a_{ij}\in\{0,1,2,$...\} of a corresponding adjacency matrix $a$ [36]. The number of outgoing and incoming links of node $i$ are denoted by $k_{i}=\sum_{j}a_{ij}$ and $q_{i}=\sum_{j}a_{ji}$ respectively, $M=\sum_{ij}a_{ij}=\sum_{i}k_{i}=\sum_{j}q_{j}=N\bar{k}=N\bar{q}$ denotes the number of directed links, and $\bar{k}$ and $\bar{q}$ the average out- and in-degree of the nodes of the network  [^1] Note that, usually, $2M=\sum_{ij}a_{ij}$ is used for nondirected networks. Here $M=\sum_{ij}a_{ij}$ is preferred since it works for both, directed and non-directed networks.. Non-directed networks are represented with symmetric adjacency matrices, in which case the number of non-directed links is $M/2$, $k_{i}=q_{i}$ and, by convention, $a_{ii}\in\{0,2,4,$...\} for each $i$.
@@ II.5 Interaction and taxation modes over networks
Different taxation and interaction modes can be considered. Each interaction-taxation event is represented by a tuple $s=(uv,l,m)$ where the pair $uv$ denotes a directed link chosen from a network $a$ representing an interaction initiated by an agent $u$ over an agent $v$, and $l$ and $m$ represent the pair of agents that are being taxed. Without loosing generality, at each event two agents are taxed simultaneously to unify different taxation modes within the same representation. The outcomes of $uv$, $l$ and $m$ affect the statistics of $b$ and $c$. Specifically, $P(c_{ij}=1|s,a)=P(c_{ij}=1|uv)=\delta_{ij,uv}+\delta_{ij,vu}$, $P(c_{ij}=0|s,a)=1-P(c_{ij}=1|s,a)$, $P(b_{i}=1|s,a)=P(b_{i}=1|l,m)=\delta_{il}+\delta_{im}$ and $P(b_{i}=0|s,a)=1-P(b_{i}=1|s,a)$.
Two interaction modes are considered. In Interaction Mode A (IMA), a directed link $uv$ is chosen uniformly at random among the $M$ available. This occurs with probability
@@
For each interaction mode, two taxation modes are also considered. In Taxation Mode A (TMA), the agents $u$ and $v$ participating in the interaction $uv$ are the ones being taxed. Formally, $l=u$ and $m=v$, so $P(l,m|uv)=P(l|uv)P(m|uv)=\delta_{lu}\delta_{mv}$. Therefore
@@
where $h_{i}=\sum_{u}a_{ui}/k_{u}\approx\sum_{u}a_{ui}/\bar{k}=q_{i}/\bar{k}$ if $1/k_{u}\approx 1/\bar{k}$ is assumed for all $u$ in the neighborhood of $i$. In Taxation Mode B, the pair of agents being taxed are uniformly chosen at random. Hence, $P(l,m|uv)=P(l|uv)P(m|uv)=N^{-2}$ and, therefore
@@ III Results
Figure 1: Behavior of EYS model on fully connected networks. a) The expected relative wealth $x_{1}$ owned by the richest agent at the stationary regime is plotted vs temperature $T$ for different system sizes $N$, averaged over 64 MC simulations of the stochastic process of Eq. II.1 starting from uniform initial conditions. b) A comparison of $x_{1}$ vs $T$ for $N=128$, showing results from MC simulations (blue circles), a single numerical simulation of the QMF approximation of Eq. II.4 starting from a uniform initial condition with added Gaussian noise (mean = 0, standard deviation = 0.01, green squares), and the MF prediction of Eq. 13 (red solid line). c) In a system of $N=16$ agents, each line represents the time evolution of the $w_{i}/W$ for i=1,,16i=1,...,16 in a MC simulation below the critical temperature, $T=0.25$. The inset shows analogous results but above the critical temperature, $T=1.1$. d) Time evolution of the same system as in c) but for a single numerical simulation of the QMF approximation at $T=0.25$ (main plot) and $T=1.1$ (inset).
@@ III.1 Fully connected networks
Figure 1a) shows results for MC simulations of the stochastic process of Eq. II.1 for different sizes $N$, which corresponds to the EYS model on fully connected networks. Except when stated otherwise, the uniform state $x_{i}=1/N$ is used as the initial condition for all simulations. The expected relative wealth of the richest agent $x_{1}$ in the steady state is plotted vs the temperature $T$. The phase transition is confirmed and finite size effects are observed. In Fig. 1b) the MC simulations are compared against numerical predictions of the QMF approximation of Eq. II.4 and the analytical MF predictions derived from Eq. III.1. Here, a small Gaussian noise is added to the initial condition to break the symmetry. As can be seen, the predictions agree up to finite size effects. Examples of the microscopic dynamics from simulations of the MC and the QMF approximations of a system with $N=16$ agents are shown in Figs. 1c) and 1d), respectively. Moreover, the wealth of multiple agents increases with time during an initial transient period. Note, however, this growth is short-lived. Eventually, all agents with the exception of the richest begin to experience a decline in wealth. This behavior arises because wealthier agents accumulate resources at the expense of their neighbors. During the early stages, several secondary wealthy agents see their wealth grow while they extract resources from poorer neighbors. Once these surrounding resources are exhausted, only the competition between the rich agents remain, ultimately favoring only the wealthiest.
@@ III.2 Erdős-Rényi random networks
The numerical experiments show that, for all interaction and taxation modes, the value of $\left\langle x_{1}\right\rangle_{a}$ grows with $\bar{k}$ but decays with $N$ at zero or small temperatures, $T\gtrsim 0$ (not shown). In other words, there is a competition between $\bar{k}$ and $N$ for consolidating or weakening the condensed phase, respectively. To better characterize this behavior, assume the scaling hypothesis
@@
for arbitrary values of $\lambda$, $\bar{k}$, $N$ and zero or small temperatures. Choosing the value of $\lambda$ such that $\lambda^{\mu}N=1$, the scaling relation
@@
is obtained, for some universal function $g_{T}$ that generally depends on $T$. As shown in Fig. 3, a collapse of the curves $N^{-1/\mu}\left\langle x_{1}\right\rangle_{a}(\bar{k},N)$ vs $N^{-\nu/\mu}\bar{k}$ is obtained for different system sizes $N$ while varying $\bar{k}$. By visual inspection, it is found that the scaling occurs for $\nu=\mu=\infty$ such that $\nu/\mu=1$ when $T=0$ (Figs. 3 a) and b)) and for $\nu=\mu=-4$ when $T=0.06$ (Figs. 3 c) and d)). In other words, the scaling
@@
Figure 3: Scaling of the network average of the expected relative wealth of the richest agent $x_{1}$ at the stationary regime, with respect to network sizes $N$ and average degrees $\bar{k}$, for MC simulations (left column) and QMF approximations (right column), at two specific temperatures, $T=0$ (top row) and $T=0.06$ (bottom row). In all cases, the stars are obtained at $\bar{k}=N-1$, i.e. the fully connected case. In panel d), the dashed cyan curve corresponds to the scaling $g_{T}(z)=(1-z)^{-1/4}$ for $z=(N-1)/N\lesssim 1$ and the red solid curve to the scaling $g_{T}(p)=-1/3+p^{1/3}(1-p)^{-1/4}$ for $p\approx\bar{k}/N$. In the inset, the same results are shown in log-log scale.
@@
which captures the imbalance between poor and rich agents. By construction, $\phi\to 0$ as $T\to\infty$, and $\phi\approx 1$ in the low-temperature and high-connectivity limit $T\to 0$, $\bar{k}\to N-1$. These predictions for $\phi$ are corroborated by both MC simulations and QMF approximations, as shown in Figs. 5 a) and b). In particular, Fig. 5a) reveals a crossover at a characteristic temperature $T=T_{c}$, which appears consistently across different values of $\bar{k}$. Moreover, while the transition is continuous in MC simulations, it becomes discontinuous in the QMF approximation as $\bar{k}\to N-1$, likely due to the mean-field treatment. The behavior of the time-averaged order parameter $\langle\phi\rangle_{a}$ also reflects the interplay between connectivity $\bar{k}$ and system size $N$. For $T<T_{c}$, $\langle\phi\rangle_{a}$ increases with $\bar{k}$ and decreases slightly with $N$; the opposite trend is observed for $T>T_{c}$. These trends are consistent with the approximation $r_{\max}\approx(2/\bar{k})\ln N$ derived from Eq. 21 for $\bar{k}\ll N$ and $T\approx 0$, which predicts a logarithmic growth with $N$ and an inverse dependence on $\bar{k}$.
@@
In Fig. 5d) network averages of the cumulative wealth of the rich agents, $u^{*}R^{*}$ are plotted as a function of $T$ for different network sizes $N$ and a fixed link density $p\approx\bar{k}/N$ for QMF approximations (main plot) and MC simulations (inset). A clear change in behavior is observed just below $T_{c}$, in agreement with the existence of a phase transition. In particular, for QMF approximations, there is a non-monotonous behavior where the minimum of $u^{*}R^{*}$ drops with $N$ because $R^{*}$ decays faster than the growth of $u^{*}$ at such point. This behavior is not observed in MC where $R^{*}$ tends to overestimate the effective number of rich agents due to stochastic fluctuations.
@@
Figure 5: a) The steady state network average of the order parameter $\phi$ defined in Eq. 22 is plotted as a function of temperature $T$ for networks of $N=128$ nodes and different average degrees $\bar{k}$, using IMA and TMA, for QMF approximations. In the inset, analogous results for for MC simulations. b) Similar to panel a) and corresponding inset, but fixing the average degree $\bar{k}=8$ while varying the network size $N$. c) The network average of the average wealth $u^{*}$ of the rich agents is plotted as a function of $T$ by direct calculation with MC simulations (orange circles) and QMF approximations (purple symbols) for $N=128$ and $\bar{k}=8$. For comparison, corresponding predictions from Eq. 19 obtained by direct calculations of network averages of the number $R^{*}$ of rich agents are included for MC simulations and QMF approximations (solid lines of respective colors). d) The network average of the aggregated wealth of the rich agents $R^{*}u^{*}$ is plotted as a function of the temperature $T$ for different values of $N$, keeping fixed the link density $\bar{k}/N=0.5$, for QMF approximations (main plot) and MC simulations (inset).
@@
In Fig. 6 the network average of the distribution of normalized wealths $P(x)$ is shown for QMF approximations in different scenarios. The top row shows distributions for all interaction and transaction modes for relatively sparse networks. In the middle and bottom row, distributions for the IMA/TMA modes are shown for varying degrees $\bar{k}$ and network sizes $N$, respectively. In the left column, the distribution is shown for the condensed phase. In all cases two modes or peaks are observed, one for poor agents and another for rich agents (Figs. 6 a)d) and g)). This result has also been observed in the MC simulations. In the middle column, distributions obtained for temperatures near or below the critical point are shown. Broad distribution are observed, except for the IMB/TMA case, where there is a sharp cutoff at the tail of the curve (Fig. 6b)). The distribution broadens as the average degree $\bar{k}$ decreases for IMA/TMA (Fig. 6e)). Interestingly, this effect is also observed at high temperatures (Fig. 6f)). However, as seen in Fig. 6c), this is not the case for the other modes, especially for TMB where the distribution is highly concentrated around $x=1/N$. Finally, as shown in Figs. 6 g)h) and i), the distributions suffer little variation with network size.